Find the slope of any line parallel to the line through points (14, 1) and (1, 2).

Answer by Nate(3500)

slope = (y-y)/(x-x)
Find the slope of any line parallel to the line through points (14, 1) and (1, 2).
slope = (1-2)/(14-1) = (-1)/(13) = -1/13